OpenStudy (anonymous):

If 9a^2-b^2=0 and ab<0, then (a-b)/(a+b)=?

OpenStudy (agent0smith):

This might help somewhat...\[\Large 9a^2-b^2= (3a)^2 - b^2\]Notice this is a difference of two squares.

OpenStudy (anonymous):

what do you mean? i cannot understand

OpenStudy (anonymous):

use x ^2 - y ^2 = (x+y)(x-y)

OpenStudy (anonymous):

i know, but how about the next step

OpenStudy (anonymous):

i think u hv got x = 3a and y = b

OpenStudy (anonymous):

yes

OpenStudy (agent0smith):

\[\Large (3a)^2 - b^2 = (3a + b)(3a - b) = 0\]

OpenStudy (anonymous):

just equate them to 0

OpenStudy (anonymous):

b=3a?

OpenStudy (anonymous):

-1/2 ?

OpenStudy (anonymous):

but ab <0 so b not = 3a

OpenStudy (anonymous):

ohh..

OpenStudy (anonymous):

b=-3a

OpenStudy (anonymous):

u hv to equate the other one 3a+b = 0

OpenStudy (anonymous):

yeah just replace b with -3a

OpenStudy (anonymous):

what did u get ??

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

gr8

OpenStudy (anonymous):

well done

OpenStudy (anonymous):

thx:)